# FALCON: An ML Framework for Fully Automated Layout-Constrained Analog Circuit Design

### v1

@@ E.1 Design Rule Enforcement in 45 nm CMOS
  • Inductor-to-inductor spacing: $\geq$ 35.0 $\mu$m to mitigate mutual inductive coupling and magnetic interference.
  • Guardring placement: Sensitive analog blocks are enclosed by N-well or deep N-well guardrings with spacing $\geq$ 5.0 $\mu$m to suppress substrate noise coupling.
  • Differential pair symmetry: Differential signal paths are layout-matched to ensure $\Delta L<0.5\,\mu$m, minimizing mismatch and preserving phase balance.
@@
MIM Capacitor (QT, LD, VV, OB) QT/LD Minimum Cap Width $W_{\text{MIN}}$ 6.05 $\mu$m
QT/LD Maximum Cap Width $W_{\text{MAX}}$ 150.0 $\mu$m
QT/LD Cap Length $L$ 6.0 $\mu$m
VV VV Square Size VV_SIZE 4.0 $\mu$m
VV VV Spacing VV_SPACE 2.0 $\mu$m
VV VV to Edge Spacing VV_EDGE_MIN 1.0 $\mu$m
Resistor (RX, CA, M1) RX Minimum Width $W_{\text{MIN}}$ 0.462 $\mu$m
RX Maximum Width $W_{\text{MAX}}$ 5.0 $\mu$m
RX Minimum Length $L_{\text{MIN}}$ 0.4 $\mu$m
RX Maximum Length $L_{\text{MAX}}$ 5.0 $\mu$m
CA Contact Size CA_SIZE 0.06 $\mu$m
CA Contact Spacing CA_SPACE 0.10 $\mu$m
CA CA to Edge Spacing CA_EDGE 0.11 $\mu$m
Inductor (M3) M3 Minimum Width M3_W_MIN 2.0 $\mu$m
M3 Maximum Width M3_W_MAX 20.0 $\mu$m
M3 Minimum Spacing M3_S_MIN 2.0 $\mu$m
Grid All Layers Minimum Grid Min_Grid 0.005 $\mu$m
@@
VV Via layer connecting different metal layers, especially in capacitor and dense routing regions.
@@ E.2 MIM Capacitor Capacitance Model
where $L$ and $W$ are the layout length and width in $\mu$m, $C_{a}$ is the area capacitance density, and $C_{p}$ is the fringing field contribution per unit length. This model includes both area and perimeter contributions to more accurately reflect layout-dependent capacitance in IC design (see Figure 14(a)).
@@ 1. Area Capacitance Term:$C_{a}\cdot L\cdot W$
2. Perimeter (Fringing) Capacitance Term:$C_{p}\cdot 2\cdot(L+W)$
@@ E.3 N+ Silicided Polysilicon Resistor Model
  • $W=5.0\,\mu$m is the drawn width; $\Delta W=0.048\,\mu$m accounts for process-induced width bias.
@@ E.4 Octagon Spiral Inductor Model
$L=2.454\times 10^{-4}\cdot D_{\text{out}}^{-1.21}\cdot W^{-0.163}\cdot D_{\text{avg}}^{2.836}\cdot S^{-0.049}\quad[$\mathrm{nH}$]$
@@
Resistor Length $L$ 4.32–20.42 $\Omega$ 10.44–34.36 $\mu\mathrm{m}^{2}$ Series if $>$ max, parallel if $<$ min
Capacitor Width $W$ 46.32–1042.4 fF 176–3344 $\mu\mathrm{m}^{2}$ Parallel if $>$ max, series if $<$ min
Inductor Radius $R$geq 0.1 nH $\geq$ 5640 $\mu\mathrm{m}^{2}$ Continuous radius scaling
@@ E.5 Layout Examples of Synthesized Circuits
To illustrate the correspondence between schematic and layout representations, we present two representative synthesized circuits: an IFVCO and a DLNA, shown in Figure 15 and Figure 16, respectively.
@@ F.2 Limitations
This work focuses on a representative set of 20 curated analog topologies spanning five circuit families. While this enables rigorous evaluation and benchmarking, extending support to arbitrary user-defined netlists is a natural next step. Notably, our GNN demonstrates strong generalization even to topologies not seen during training (e.g., RVCO), suggesting broad applicability within this space.
All interconnect routing in the layout phase is currently performed manually, as the flow does not yet integrate a full analog router. This decision allows precise control over parasitic management and DRC compliance but limits scalability for more complex designs. Integration with automated layout tools can further streamline this process.
We do not perform repeated training runs for each model configuration, as the pipeline is designed to be computationally efficient and executable on CPU-only hardware. Unlike benchmarking-focused pipelines, our flow incorporates analog simulation and layout-aware inference, both of which are costly to rerun at scale. Moreover, simulation and layout processes are deterministic, eliminating the need for result averaging over multiple random seeds.

### v2

@@ E.1 Design Rule Enforcement in 45 nm CMOS
[…]
@@
[…]
@@
$V$ Via layer connecting different metal layers, especially in capacitor and dense routing regions.
@@ E.2 MIM Capacitor Capacitance Model
$C_{N}=C_{a}\cdot L\cdot W+C_{p}\cdot 2\cdot(L+W)\quad\text{[fF]}$
@@
[…]
@@ E.3 N+ Silicided Polysilicon Resistor Model
[…]
@@ E.4 Octagon Spiral Inductor Model
$L=2.454\times 10^{-4}\cdot D_{\text{out}}^{-1.21}\cdot W^{-0.163}\cdot D_{\text{avg}}^{2.836}\cdot S^{-0.049}\quad[$\mathrm{nH}$\$]
@@
Resistor Length $L$ 4.32–20.42 $\Omega$ 10.44–34.36 $\mu\mathrm{m}^{2}$ Series if >> max, parallel if << min
Capacitor Width $W$ 46.32–1042.4 fF 176–3344 $\mu\mathrm{m}^{2}$ Parallel if >> max, series if << min
Inductor Radius $R$ $\geq$ 0.1 nH $\geq$ 5640 $\mu\mathrm{m}^{2}$ Continuous radius scaling
@@ E.5 Layout Examples of Synthesized Circuits
To illustrate the correspondence between schematic and layout representations, we present three synthesized circuits: DBAMixer, IFVCO, and DLNA, shown in Figures 15, 16, and 17, respectively.
(a) Designed DBAMixer schematic
(b) Layout of designed DBAMixer
Figure 15: Stage 3 results for a synthesized DBAMixer. The schematic (a) reflects optimized parameters to meet the target specification. The layout (b) is DRC-compliant and physically realizable. The final design achieves a mean relative error of 0.2% compared to the target performance.
@@ F.2 Limitations
This work focuses on a representative set of 20 curated analog topologies spanning five circuit families. Consequently, the topology selection stage is limited to suggesting only among the designs present in the training set and cannot synthesize novel circuits. A natural future direction is to either extend the training library to a broader set of topologies or replace the classifier with a generative model capable of directly proposing new netlists conditioned on input specifications. In contrast, the GNN-based forward modeling stage is designed to operate on arbitrary circuit graphs and has already demonstrated strong generalization to unseen architectures (e.g., RVCO), indicating that no modification to this stage is required to support novel circuits.
Beyond topology considerations, the dataset is constructed at a fixed operating frequency of 30 GHz, which ensures consistency across circuit families but constrains frequency generalization. Although the framework can, in principle, extend to other operating points—for example, the voltage amplifier topologies already demonstrate scalability across varying gain–bandwidth trade-offs—systematic validation across diverse frequency bands is beyond the scope of this work. Extending the dataset to cover multiple operating frequencies, or incorporating frequency as an explicit conditioning variable during training, represents an important direction for broadening applicability.
Finally, the differentiable layout model in FALCON captures parasitic effects through analytical approximations of passive components, which is effective for guiding parameter optimization within the learning framework. However, this approach does not fully replace electromagnetic (EM) simulations or post-layout verification, and electromigration constraints are not explicitly incorporated. Incorporating learned parasitic estimators, EM-informed models, and reliability constraints, therefore, remains an important extension toward bridging schematic-level optimization and silicon-proven robustness. In addition, all interconnect routing in the current flow is performed manually to ensure precise control over parasitic management and DRC compliance. While this provides accuracy for the studied designs, it limits scalability for more complex circuits, motivating future integration with automated analog routing tools.